OpenStudy (anonymous):

what is the formula used to get the area of a triangle knowing the measure of its sides a, b, c?

OpenStudy (anonymous):

you can use "heron's formula"

OpenStudy (anonymous):

area = sqrt(s(s-a)(s-b)(s-c) if my memory serves me right

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Heron%27s_formula

OpenStudy (anonymous):

jimmyrep has it. of course you need to know what "s" is

OpenStudy (anonymous):

where s = sum of the 3 sides / 2

OpenStudy (anonymous):

First find p, the perimeter of the triangle by adding all three sides. Then use sqrt(p(p-a)(p-b)(p-c)) where a, b, and c are sidelengths.

OpenStudy (anonymous):

"semiperimeter" good word that, even if spell check doesn't like it

OpenStudy (anonymous):

@smoothmath it is not p, it is what jimmyrep said

OpenStudy (anonymous):

ok ok then a complete equation would be?

OpenStudy (anonymous):

jimmyrep has it or look at the wiki entry. it has it all written out

OpenStudy (anonymous):

\[A = \sqrt{s(s - a)(s - b)(s - c)},\]where s = (a + b + c)/2 [Heron's formula]

OpenStudy (anonymous):

area = sqrt[(s(s-a)(s-b)(s-c)]

OpenStudy (anonymous):

yep abtrehearn has it too

OpenStudy (anonymous):

you can also find one of the angles using the law of cosine and then use \[Area=\frac{1}{2}ab\sin(C)\]

OpenStudy (anonymous):

thank you all, now I would like to know how would it look like in 3 dimensions adding z to the formula, that is volume. Then how can I calculate the weight of that 3d object by knowing the average density of it?

OpenStudy (anonymous):

If you rewrite the formula without the s and express only in terms of a, b, and c, then \[A = (1/4)\sqrt{(a + b + c)(-a + b + c)(a - b + c)(a + b + c)}.\]

OpenStudy (anonymous):

\[(1/4)\sqrt{(a + b + c)(-a + b + c)(a - b + c)(a - b - c)}.\]

OpenStudy (anonymous):

but now I have a,b,c,d,e,f, its a pyramid

OpenStudy (anonymous):

ooh I found it, in wikipedia 5.1 Heron-looking formula for tetrahedra